OpenStudy (anonymous):

Jerome walked 2/3 mile in the morning and 11/12 mile in the evening. How many miles did he walk in all?

OpenStudy (anonymous):

2/3 + 11/12 = 8/12 + 11/12 = 19/12 = 1 7/12 miles